1. A construction company has bid on two contracts. The probability of winning contract A is 0.3. If the company wins contract A, then the probability of winning contract B is 0.4. If the company loses, contract A, then the probability of winning contract B decreases to 0.2. Find the probability of the following events:(full details)
A. Winning both contracts.
B. Winning exactly one contract.
C. Winning at least one contract.

Answer by Boreal(15235)

winning both is 0.3*0.4=0.12
winning neither 0.7*0.8, the complement of winning=0.56
winning at least 1 is the complement of the second one, or 0.56, and that is 1-0.56=0.44
winning A alone is 0.3*0.6=0.18
winning B alone is 0.7*0.2=0.14
Ignoring the at least 1, the other four add to 1.0, which they must.